OpenStudy (anonymous):

Find the value of y so that the line passing through (6, y) and (7, 5) has a slope of -3.<--help please!

OpenStudy (anonymous):

Slope = Change of Y / Change of X ; To find change of Y you take the difference between Y values and to find change of X you take the difference of X values. For example, if you had (X1,Y1) and (X2,Y2) the slope would be (Y2-Y1)/(X2-X1) or (Y1-Y2)/(X1-X2). Both are the same slope. So slope= (Y2-Y1)/(X2-X1) [I chose this one but you can use the other if you want to). Now solve. Slope = -3. You have values for X1,X2,Y1, solve for Y2

OpenStudy (anonymous):

what do i put in for y1 in the equation since it is just y?

OpenStudy (across):

You have to solve the following equation:\[\frac{5-y}{7-6}=-3.\]

OpenStudy (anonymous):

i got 5y/1=-3 u could change the 5y/1 to just 5y so you have 5y=-3 is that right?

OpenStudy (across):

Where did you get that 5y from?

OpenStudy (anonymous):

No it is 5-y/1 =-3. Not 5y. So it becomes 5-y=-3.

OpenStudy (anonymous):

ohh yea i accidently multiplied the 5 and y my bad

OpenStudy (anonymous):

i solved it y=2 right?

OpenStudy (anonymous):

*its

OpenStudy (across):

Almost:\[5-y=-3,\]\[y=8.\]

OpenStudy (anonymous):

Don't forget a negative minus a number is negative. So -3 minus 5 will become -8. So then -y=-8, which means you divide by negative one so you get y=8

OpenStudy (anonymous):

dont you add 5 since you have a negative so you do the opposite which wud be addition?

OpenStudy (anonymous):

To isolate the system you subtract the 5. 5-y=-3. subtract 5 from both sides to get rid of the 5 on the left side. you get -y=-8

OpenStudy (anonymous):

ohh i see thats always confusing to me thank you both

OpenStudy (anonymous):

No problems. Good luck ^_^